OpenStudy (anonymous):

integrate x^3/sqroot 16-x^2 dx

OpenStudy (dumbcow):

u = 16- x^2 du = -2x dx \[=-\frac{1}{2}\int\limits_{}^{}\frac{16-u}{\sqrt{u}}du\]

OpenStudy (dumbcow):

can you do it from there

OpenStudy (anonymous):

i am confuse